Determine whether or not the following series converges. ∑_(k=1)^∞▒8k/√(k^2+5)


1
Expert's answer
2020-11-09T16:12:17-0500

"∑_{k=1}^∞\frac{8k}{\sqrt{(k^2+5)}}"

Keeping only the fastest growing terms in the numerator abd denominator, we have;

"∑_{k=1}^∞\frac{8k}{\sqrt{k^2}}=∑_{k=1}^∞\frac{8k}{k}=\\
∑_{k=1}^∞8"

p = 0

This is a p-series with p "\le" 0 so it is divergent.


We use the limit comparison test to compare the series.

"∑_{k=1}^∞ak = ∑_{k=1}^∞\frac{8k}{\sqrt{(k^2+5)}}"

"\u2211_{k=1}^\u221ebk = \u2211_{k=1}^\u221e8"


"\begin{aligned}
\therefore \dfrac{ak}{bk} &= \dfrac{\frac{8k}{\sqrt{(k^2+5)}}}{8}\\
\\
&= \dfrac{16k}{\sqrt{(k^2+5)}}\\
\\
&= \dfrac{\sqrt{256k^2}}{\sqrt{(k^2+5)}}\\
\\
&= \sqrt{\dfrac{256k^2}{(k^2+5)}}\\
\\
&= \sqrt{\dfrac{1}{1/256 + 5/k^2}}\\
\\
&\to 1= L
\end{aligned}"


According to the limit comparison test, if 0<L<"\infin", we can conclude that both series are divergent if one of them is.


"\therefore ∑_{k=1}^∞\frac{8k}{\sqrt{(k^2+5)}}" is divergent